CH 24 #15

Assume the mostly infrared radiation from a heat lamp acts like a continuous wave with wavelength 1.50 µm.

(a) If the lamp's 205 W output is focused on a person's shoulder, over a circular area 26.0 cm in diameter, what is the intensity in W/m2?

(b) What is the peak electric field strength in kV/m?

(c) Find the peak magnetic field strength in µT.

(d) How long will it take in seconds to increase the temperature of the 4.10 kg shoulder by 2.00°C, assuming no other heat transfer and given its specific heat is 3.47 ✕ 103 J/(kg · °C)? seconds

Pleas show step by step with legible writing. Thank you
Answer #1

a)

Intensity is given by

I=P/A =205/[pi*(0.26/2)2]

I=3861.16 W/m2

b)

SInce

E=1705 V/m =1.705 KV/m

c)

B=E/c =1705/(3*108)

B=5.6835*10-6 T =5.6835 uT

d)

Heat generated

Q=mc(T2-T1)=4.1*(3.47*103)*2=28454 J

Therefore

t=Q/P =28454/205

t=138.8 s